In Ms. Pham's social studies class, a test is worth four times as much as homework. If a student has an average of 85% on tests and 95% on homework, what is the student's average?
Answer by flame8855(424)

0.85 * 4 = 3.4
3.4 + 0.95 = 4.35
4.35/5=0.87
87 %